# Sheshathanks: Shesha or Sheshanaga is a serpentine demigod in Hindu mythology with multiple heads and serves as the celestial bed of Lord Vishnu.  : Multi-head Microarchitectural Leakage Discovery in new-generation Intel Processors

### v1

@@ 8.3 Attack on Cryptographic Applications
To test the extent of FMA leakage in AMM/AMS, we implement the MulALPart algorithm from [17] as depicted in Listing 2 (MulAHPart follows suit, but uses VPMADD52HUQ instead of VPMADD52LUQ). Similar to Sec. 8.2.2, the inputs to AMM/AMS are programmed to 0xfa. MulALPart performs efficient left-shift in AMM/AMS implementation [17], such that this result can directly be used in other IFMA friendly operations in AMM/AMS. Note that AMM/AMS can be performed on any arbitrary integer $A$ exceeding the width of FMA registers, thereby requiring a total of $z$ logical registers (notated as $\{A_{1},A_{2},...,A_{z}\}$). Likewise, $i$ represents the index of the current chunk to be operated. $A_{curr}$ is then assigned the broadcasted value of $A[i]$, while $\{X_{1},X_{2},...,X_{i},...,X_{z}\}$ are temporary scratch variables. First, as in Line $4$, $X_{i}$ is assigned the value VPMADD52LUQ($X_{i},A_{curr},A_{i}$). Then, in a tight loop from $i+1$ to $z$, the temporary products VPMADD52LUQ(ZERO, $A_{curr}$, $A_{i}$) are computed, left-shifted, and added to previously computed $X_{j}$. From our discussion in Sec. 8.2, it is clear that the third input operand to IFMA instructions is vulnerable to leakage. With GCC, it is $A_{i}$ that is emitted to be the third operand. This is intuitive: note that in VPMADD52LUQ(ZERO, $A_{curr}$, $A_{i}$), the first two operands are fixed for the entire duration of the loop and thus are pre-loaded by the compiler (as first and second operands of VPMADD52LUQ) in registers zmm1 and zmm3. Since $A_{i}$ changes in every iteration, this becomes the third (memory) operand of VPMADD52LUQ. Therefore, the leaked value is the multiplicand ${A_{i}:i\in\{i+1,i+2,...,z\}}$, which constitutes the actual input $A$ to AMM/AMS.
@@ 8.3.3 Supersingular Isogeny Key Encapsulation
In [13], the authors use IFMA to accelerate SIKE, which is a post-quantum isogeny based cryptosystem (similar to SIDH in Sec. 8.3.2). Here too, IFMA is used to speed upbase/extension field arithmetic, point arithmetic, and isogeny computations. Since SIKE relies mainly on SIDH-friendly primes, techniques for speeding up modular reduction also apply here. Henceforth, as with Sec. 8.3.2, this implementation [^9] https://gitlab.uni.lu/APSIA/AVXSIKE is also vulnerable to the leakage demonstrated here.
@@ 8.3.4 Other Use-Cases
We briefly mention other use-cases which have similar vulnerability. In [46], the authors propose IFMA based Montgomery multiplication, which again leaks the operands through the leakage described here. Likewise, in [59], Montgomery reduction is also sped up using IFMA, and is vulnerable to our attack. Note that [59] speeds up Dilithium, which is a post-quantum digital signature scheme, and has similar vulnerability as AMM in Sec. 8.3.1 of leaking the plaintext message being signed. Finally, recent works like [7] have also demonstrated the use of IFMA to speed up homomorphic operations as well, and is also vulnerable to the leak [^10] It is unclear, though, what the extent of exploitability is in the case of homomorphic encryption. This is because even though there is leakage, all data is essentially encrypted and thus secure by the underlying guarantees of the homomorphic encryption scheme..
@@ 10 Mitigations
Precision based assists and machine clears: As pointed out in Sec. 6.2, the SMC occuring due to the precision during hardware assist is managed as loads/stores. Thereby, an lfence before and after an instruction sequence with a homogeneous precision type (either double or single precision) serializes the load and eliminates transient execution.
SIMD-Vector Transition based SMC: As discussed in Sec. 6.1, SIMD-Vector transitions based SMC occur to effectively issue blend operations. As Intel Optimization Manual suggests [23], SIMD-Vector based SMC can be prevented by appropriate use of zero-latency instructions vzeroupper/vzeroall before and after a function executing SIMD or Vector instruction sequences.
@@ 11 Conclusion
Appendix A Precision based Transient Execution Paths
A.1 Precision based Hardware Assist
A hardware assist (represented by ASSISTS.HARDWARE in Tab. 1) is required when an exceptional scenario arises where dedicated hardware extensions are needed to execute the concerned instruction. For performance optimizations, Intel provides dedicated hardware for a variety of operations. To exemplify a few, there exist AES-NI execution units for cryptographic operations, hardware support for half-precision conversion operations using VCVTPS2PH, and so on. While the programmer can interface with these hardware extensions separately through ISE instructions, a dedicated hardware assist is needed when an exceptional situation arises where intervention of such hardware is required. One such scenario unearthed by Shesha is the requirement of hardware assists in converting representations of floating point numbers. Without loss of generality, consider DP xmm0, xmm1, xmm2 followed by SP xmm3, xmm4, xmm0. This instruction sequence updates xmm0 to contain a double-precision floating-point number. However, without any explicit conversion instruction in between, the subsequent instruction SP xmm3, xmm4, xmm0 is forced to rely upon a hardware assist to convert xmm0’s natural double-precision floating point representation to a single-point representation, thereby creating a transient execution window.
A.2 Precision based SMC Machine Clear
As evident from Sec. A.1, mismatches in the precision of a register involved in a RAW data dependency incur hardware assists for fixing the mismatching representations. However, considering DP xmm0, xmm1, xmm2 followed by SP xmm3, xmm4, xmm0, we note that without the presence of an explicit conversion operation (like CVTPD2PS or CVTPS2PD) between the double-precision instruction DP and the single-precision instruction SP, the processor is faced with two challenges. First, it needs a way to represent xmm0 as single-precision floating-point value consumable by instruction SP (already detailed in Sec. A.1). Secondly, for architectural correctness, it needs a way to retain the initial value of xmm0 in double-precision at the end of execution of DP, even though a corresponding single-precision representation is needed for SP.
Note that architecturally, instruction sequence DP xmm0, xmm1, xmm2 followed by SP xmm3, xmm4, xmm0 requires xmm0 to be still double precision after SP ends execution. Hence, architecturally, this code sequence must behave as if two implicit conversion instructions were written into the code sequence by the programmer- DP xmm0, xmm1, xmm2, followed by an explicit single precision conversion CVTPD2PS xmm0, xmm0, followed by the single precision SP xmm3, xmm4, xmm0, and then followed by an explicit conversion back to double precision CVTPS2PD xmm0, xmm0. Since the programmer has not included these explicit conversions, the hardware assist described in Sec. A.1 cannot guarantee architectural correctness without auxiliary code to manage xmm0’s initial double precision value, and restoring it later. This management code is exactly what the SMC machine clear is performing in this context. To ensure this further, we rely upon instruction cache metrics ICACHE.ACCESSES and ICACHE.MISSES, counting the number of total accesses to the instruction cache and corresponding misses respectively. Averaged across 1 million executions, we observe ICACHE.ACCESSES to be $21$ for a SP-SP or a DP-DP instruction sequence, while it climbs to $36$ for SP-DP or DP-SP instruction sequences. Moreover, we observe no ICACHE.MISSES even though (on average) $15$ more instructions are added to the instruction cache as a result of the SMC. This reinforces the idea that all instructions needed to manage xmm0’s double precision state while executing a single precision operation are directly added to the instruction cache. And these newly added instructions force re-steering the pipeline, squashing any in-flight instructions, and re-issuing these instructions. This becomes the root cause of the precision-based SMC machine clear, thereby creating yet another novel transient execution path.
A.3 Precision based Memory Ordering
In Sec. A.1 and Sec. A.2, we detailed how two separate and novel transient executions paths are uncovered by Shesha in order to manage the state of a SIMD register upon implicit precision conversions. We note that simply from observing an SMC machine clear (cf. Sec. A.2), it is not possible to uncover what exact instructions are being added to the instruction cache. However, yet another transient execution path uncovered by Shesha offers more insights into the exact nature of this SMC - occurrence of memory ordering issues during SP-DP or DP-SP without any explicit conversions. We delve deeper into such memory ordering in this section.
Again, without loss of generality, we consider the instruction sequence DP xmm0, xmm1, xmm2 followed by SP xmm3, xmm4, xmm0 incurring a memory ordering machine clear. Concretely, since both SP and DP are exclusively arithmetic instructions, the only possibility of a memory ordering issue arises if the SMC is relying upon load/store operations to manage the initial and final state of xmm0 (cf. Sec. A.2). This hypothesis is further concretized by the fact that memory ordering issues in Intel CPU occur due to two reasons as stated in [25]- ① a load instruction does not conform to x86 memory ordering rules or ② initiation of MESI snoop cache coherence protocol. In this case, our observation and experiment design indicates no MESI initiations occur, thereby leaving the first case as the root-cause of memory ordering issue. To elaborate further, we observe zero snoop requests through measuring performance counters like MEM_LOAD_L3_HIT_RETIRED.XSNP_HIT, MEM_LOAD_L3_HIT_RETIRED.XSNP_NO_FWD, OCR.DEMAND_DATA_RD.L3_HIT.SNOOP_HITM, and other counters involving snoop requests [25], leaving only the possibility of a faulty load as the potential cause of the memory ordering issue. Moreover, should we insert an explicit lfence to obtain the sequence: DP xmm0, xmm1, xmm2, followed by lfence and then followed by SP xmm3, xmm4, xmm0, the memory ordering issues vanishes. Combining both these observations (① absence of MESI snoop requests and ② serialized load executions using lfence preventing memory ordering issues), we conclusively state that the underlying SMC to manage internal state of xmm0 utilizes at least one store/load pair, where the load is speculatively issued before the store causing a memory ordering issue. We stress that the offending store/load pair is on the same effective virtual address because we observe no memory disambiguation issues (through MACHINE_CLEARS.DISAMBIGUATION in Tab. 1). This is because memory disambiguation issues arise due to a store/load pair on actual differing effective virtual addresses being speculated to operate upon the same effective virtual address during the address resolution process.
Appendix B SIMD-Vector Transition based SMC
B.2 Speculative transition penalty in Alder Lake and Sapphire Rapids
Furthermore, our analysis shows that the SMC machine clear occurs irrespective of ① whether the ymm registers are in Modified/Unsaved state (i.e. the upper $128$ bits of the $256$-bit ymm register) are clean, as well as ② whether there is actually any register dependency upon SIMD-Vector transitions. This cements the fact that the SIMD-Vector transition SMC is indeed a pessimistic transition penalty- penalty incurred irrespective of whether there is register dependency or whether ymm register is indeed unclean. Interestingly, incurring transition penalty without actually checking for these conditions is also done by AMD in its recent EPYC processors, where spilling occurs on the higher $128$ bits of all ymm registers, irrespective of which ymm register is involved in the transition (refer Sec.2.11.62.11.62.11.62.11.6 in [2]).
Appendix C SMT Covert Channel Sender/Receiver
Algorithm 1 SMT Covert Channel sender
1:procedure sender($\mathcal{C}$)
2:    Pin to logical core $i$ on physical core $k$
3:    Sample bit b $\leftarrow\{0,1\}$
4:    if b = 1 then
6:         Trigger execution of code sequence $\mathcal{C}$
7:    else
8:         Trigger execution of code sequence $\mathcal{C}$
Algorithm 2 SMT Covert Channel receiver
1:procedure receiver($\mathcal{C}$)
2:    Pin to logical core $j$ on physical core $k$
3:    start_time $\leftarrow$ rdtscp()
4:    Trigger execution of code sequence $\mathcal{C}$
5:    end_time $\leftarrow$ rdtscp()
6:    if end_time - start_time > threshold then
7:         received_bit = $1$
8:    else
9:         received_bit = 00
Appendix D List of leaking Instructions
Listing 5: FMA/IFMA instructions that leak.
1 ; FMA instructions
2VFMADD132PD VFMADD132PS VFMADD132SD VFMADD132SS
3VFMADD213PD VFMADD213PS VFMADD213SD VFMADD213SS
4VFMADD231PD VFMADD231PS VFMADD231SD VFMADD231SS
5VFMADDSUB132PD VFMADDSUB132PS VFMADDSUB213PD
6VFMADDSUB213PS VFMADDSUB231PD VFMADDSUB231PS
7VFMSUB132PD VFMSUB132PS VFMSUB132SD VFMSUB132SS
8VFMSUB213PD VFMSUB213PS VFMSUB213SD VFMSUB213SS
9VFMSUB231PD VFMSUB231PS VFMSUB231SD VFMSUB231SS
10VFMSUBADD132PD VFMSUBADD132PS VFMSUBADD213PD
11VFMSUBADD213PS VFMSUBADD231PD VFMSUBADD231PS
12VFNMADD132PD VFNMADD132PS VFNMADD132SD
13VFNMADD132SS VFNMADD213PD VFNMADD213PS
14VFNMADD213SD VFNMADD213SS VFNMADD231PD
15VFNMADD231PS VFNMADD231SD VFNMADD231SS
16VFNMSUB132PD VFNMSUB132PS VFNMSUB132SD
17VFNMSUB132SS VFNMSUB213PD VFNMSUB213PS
18VFNMSUB213SD VFNMSUB213SS VFNMSUB231PD
19VFNMSUB231PS VFNMSUB231SD VFNMSUB231SS
20
21 ; IFMA instructions
22VPMADD52HUQ VPMADD52LUQ
Appendix E Shesha: An algorithmic perspective

### v2

@@ 8.3 Attack on Cryptographic Applications
To test the extent of FMA leakage in AMM/AMS, we implement the MulALPart algorithm from [17] as depicted in Listing 2 (MulAHPart follows suit, but uses VPMADD52HUQ instead of VPMADD52LUQ). Similar to Section 8.2.2, the inputs to AMM/AMS are programmed to 0xfa. MulALPart performs efficient left-shift in AMM/AMS implementation [17], such that this result can directly be used in other IFMA friendly operations in AMM/AMS. Note that AMM/AMS can be performed on any arbitrary integer $A$ exceeding the width of FMA registers, thereby requiring a total of $z$ logical registers (notated as $\{A_{1},A_{2},...,A_{z}\}$). Likewise, $i$ represents the index of the current chunk to be operated. $A_{curr}$ is then assigned the broadcasted value of $A[i]$, while $\{X_{1},X_{2},...,X_{i},...,X_{z}\}$ are temporary scratch variables. First, as in Line $4$, $X_{i}$ is assigned the value VPMADD52LUQ($X_{i},A_{curr},A_{i}$). Then, in a tight loop from $i+1$ to $z$, the temporary products VPMADD52LUQ(ZERO, $A_{curr}$, $A_{i}$) are computed, left-shifted, and added to previously computed $X_{j}$. From our discussion in Section 8.2, it is clear that the third input operand to IFMA instructions is vulnerable to leakage. With GCC, it is $A_{i}$ that is emitted to be the third operand. This is intuitive: note that in VPMADD52LUQ(ZERO, $A_{curr}$, $A_{i}$), the first two operands are fixed for the entire duration of the loop and thus are pre-loaded by the compiler (as first and second operands of VPMADD52LUQ) in registers zmm1 and zmm3. Since $A_{i}$ changes in every iteration, this becomes the third (memory) operand of VPMADD52LUQ. Therefore, the leaked value is the multiplicand ${A_{i}:i\in\{i+1,i+2,...,z\}}$, which constitutes the actual input $A$ to AMM/AMS.
@@ 8.3.3 Supersingular Isogeny Key Encapsulation
In [13], the authors use IFMA to accelerate SIKE, which is a post-quantum isogeny based cryptosystem (similar to SIDH in Section 8.3.2). Here too, IFMA is used to speed upbase/extension field arithmetic, point arithmetic, and isogeny computations. Since SIKE relies mainly on SIDH-friendly primes, techniques for speeding up modular reduction also apply here. Henceforth, as with Section 8.3.2, this implementation [^9] https://gitlab.uni.lu/APSIA/AVXSIKE is also vulnerable to the leakage demonstrated here.
@@ 8.3.4 Other Use-Cases
We briefly mention other use-cases which have similar vulnerability. In [47], the authors propose IFMA based Montgomery multiplication, which again leaks the operands through the leakage described here. Likewise, in [60], Montgomery reduction is also sped up using IFMA, and is vulnerable to our attack. Note that [60] speeds up Dilithium, which is a post-quantum digital signature scheme, and has similar vulnerability as AMM in Section 8.3.1 of leaking the plaintext message being signed. Finally, recent works like [7] have also demonstrated the use of IFMA to speed up homomorphic operations as well, and is also vulnerable to the leak [^10] It is unclear, though, what the extent of exploitability is in the case of homomorphic encryption. This is because even though there is leakage, all data is essentially encrypted and thus secure by the underlying guarantees of the homomorphic encryption scheme..
@@ 10 Mitigations
Precision based assists and machine clears: As pointed out in Section 6.2, the SMC occuring due to the precision during hardware assist is managed as loads/stores. Thereby, an lfence before and after an instruction sequence with a homogeneous precision type (either double or single precision) serializes the load and eliminates transient execution.
SIMD-Vector Transition based SMC: As discussed in Section 6.1, SIMD-Vector transitions based SMC occur to effectively issue blend operations. As Intel Optimization Manual suggests [24], SIMD-Vector based SMC can be prevented by appropriate use of zero-latency instructions vzeroupper/vzeroall before and after a function executing SIMD or Vector instruction sequences.
@@ 11 Conclusion
Appendix A Shesha: An algorithmic perspective
@@
Appendix B SIMD-Vector Transition based SMC
B.2 Speculative transition penalty in Alder Lake and Sapphire Rapids
Furthermore, our analysis shows that the SMC machine clear occurs irrespective of ① whether the ymm registers are in Modified/Unsaved state (i.e. the upper $128$ bits of the $256$-bit ymm register) are clean, as well as ② whether there is actually any register dependency upon SIMD-Vector transitions. This cements the fact that the SIMD-Vector transition SMC is indeed a pessimistic transition penalty- penalty incurred irrespective of whether there is register dependency or whether ymm register is indeed unclean. Interestingly, incurring transition penalty without actually checking for these conditions is also done by AMD in its recent EPYC processors, where spilling occurs on the higher $128$ bits of all ymm registers, irrespective of which ymm register is involved in the transition (refer Section $2.11.6$ in [2]).
Appendix C Precision based Transient Execution Paths
C.1 Precision based Hardware Assist
A hardware assist (represented by ASSISTS.HARDWARE in Table 1) is required when an exceptional scenario arises where dedicated hardware extensions are needed to execute the concerned instruction. For performance optimizations, Intel provides dedicated hardware for a variety of operations. To exemplify a few, there exist AES-NI execution units for cryptographic operations, hardware support for half-precision conversion operations using VCVTPS2PH, and so on. While the programmer can interface with these hardware extensions separately through ISE instructions, a dedicated hardware assist is needed when an exceptional situation arises where intervention of such hardware is required. One such scenario unearthed by Shesha is the requirement of hardware assists in converting representations of floating point numbers. Without loss of generality, consider DP xmm0, xmm1, xmm2 followed by SP xmm3, xmm4, xmm0. This instruction sequence updates xmm0 to contain a double-precision floating-point number. However, without any explicit conversion instruction in between, the subsequent instruction SP xmm3, xmm4, xmm0 is forced to rely upon a hardware assist to convert xmm0’s natural double-precision floating point representation to a single-point representation, thereby creating a transient execution window.
C.2 Precision based SMC Machine Clear
As evident from Section C.1, mismatches in the precision of a register involved in a RAW data dependency incur hardware assists for fixing the mismatching representations. However, considering DP xmm0, xmm1, xmm2 followed by SP xmm3, xmm4, xmm0, we note that without the presence of an explicit conversion operation (like CVTPD2PS or CVTPS2PD) between the double-precision instruction DP and the single-precision instruction SP, the processor is faced with two challenges. First, it needs a way to represent xmm0 as single-precision floating-point value consumable by instruction SP (already detailed in Section C.1). Secondly, for architectural correctness, it needs a way to retain the initial value of xmm0 in double-precision at the end of execution of DP, even though a corresponding single-precision representation is needed for SP.
Note that architecturally, instruction sequence DP xmm0, xmm1, xmm2 followed by SP xmm3, xmm4, xmm0 requires xmm0 to be still double precision after SP ends execution. Hence, architecturally, this code sequence must behave as if two implicit conversion instructions were written into the code sequence by the programmer- DP xmm0, xmm1, xmm2, followed by an explicit single precision conversion CVTPD2PS xmm0, xmm0, followed by the single precision SP xmm3, xmm4, xmm0, and then followed by an explicit conversion back to double precision CVTPS2PD xmm0, xmm0. Since the programmer has not included these explicit conversions, the hardware assist described in Section C.1 cannot guarantee architectural correctness without auxiliary code to manage xmm0’s initial double precision value, and restoring it later. This management code is exactly what the SMC machine clear is performing in this context. To ensure this further, we rely upon instruction cache metrics ICACHE.ACCESSES and ICACHE.MISSES, counting the number of total accesses to the instruction cache and corresponding misses respectively. Averaged across 1 million executions, we observe ICACHE.ACCESSES to be $21$ for a SP-SP or a DP-DP instruction sequence, while it climbs to $36$ for SP-DP or DP-SP instruction sequences. Moreover, we observe no ICACHE.MISSES even though (on average) $15$ more instructions are added to the instruction cache as a result of the SMC. This reinforces the idea that all instructions needed to manage xmm0’s double precision state while executing a single precision operation are directly added to the instruction cache. And these newly added instructions force re-steering the pipeline, squashing any in-flight instructions, and re-issuing these instructions. This becomes the root cause of the precision-based SMC machine clear, thereby creating yet another novel transient execution path.
C.3 Precision based Memory Ordering
In Section C.1 and Section C.2, we detailed how two separate and novel transient executions paths are uncovered by Shesha in order to manage the state of a SIMD register upon implicit precision conversions. We note that simply from observing an SMC machine clear (cf. Section C.2), it is not possible to uncover what exact instructions are being added to the instruction cache. However, yet another transient execution path uncovered by Shesha offers more insights into the exact nature of this SMC - occurrence of memory ordering issues during SP-DP or DP-SP without any explicit conversions. We delve deeper into such memory ordering in this section.
Again, without loss of generality, we consider the instruction sequence DP xmm0, xmm1, xmm2 followed by SP xmm3, xmm4, xmm0 incurring a memory ordering machine clear. Concretely, since both SP and DP are exclusively arithmetic instructions, the only possibility of a memory ordering issue arises if the SMC is relying upon load/store operations to manage the initial and final state of xmm0 (cf. Section C.2). This hypothesis is further concretized by the fact that memory ordering issues in Intel CPU occur due to two reasons as stated in [25]- ① a load instruction does not conform to x86 memory ordering rules or ② initiation of MESI snoop cache coherence protocol. In this case, our observation and experiment design indicates no MESI initiations occur, thereby leaving the first case as the root-cause of memory ordering issue. To elaborate further, we observe zero snoop requests through measuring performance counters like MEM_LOAD_L3_HIT_RETIRED.XSNP_HIT, MEM_LOAD_L3_HIT_RETIRED.XSNP_NO_FWD, OCR.DEMAND_DATA_RD.L3_HIT.SNOOP_HITM, and other counters involving snoop requests [25], leaving only the possibility of a faulty load as the potential cause of the memory ordering issue. Moreover, should we insert an explicit lfence to obtain the sequence: DP xmm0, xmm1, xmm2, followed by lfence and then followed by SP xmm3, xmm4, xmm0, the memory ordering issues vanishes. Combining both these observations (① absence of MESI snoop requests and ② serialized load executions using lfence preventing memory ordering issues), we conclusively state that the underlying SMC to manage internal state of xmm0 utilizes at least one store/load pair, where the load is speculatively issued before the store causing a memory ordering issue. We stress that the offending store/load pair is on the same effective virtual address because we observe no memory disambiguation issues (through MACHINE_CLEARS.DISAMBIGUATION in Table 1). This is because memory disambiguation issues arise due to a store/load pair on actual differing effective virtual addresses being speculated to operate upon the same effective virtual address during the address resolution process.
Appendix D SMT Covert Channel Sender/Receiver
Algorithm 3 SMT Covert Channel sender
1:procedure sender($\mathcal{C}$)
2:    Pin to logical core $i$ on physical core $k$
3:    Sample bit b $\leftarrow\{0,1\}$
4:    if b = 1 then
6:         Trigger execution of code sequence $\mathcal{C}$
7:    else
8:         Trigger execution of code sequence $\mathcal{C}$
Algorithm 4 SMT Covert Channel receiver
1:procedure receiver($\mathcal{C}$)
2:    Pin to logical core $j$ on physical core $k$
3:    start_time $\leftarrow$ rdtscp()
4:    Trigger execution of code sequence $\mathcal{C}$
5:    end_time $\leftarrow$ rdtscp()
6:    if end_time - start_time > threshold then
7:         received_bit = $1$
8:    else
9:         received_bit = 00
Appendix E List of leaking Instructions
Listing 5: FMA/IFMA instructions that leak.
1 ; FMA instructions
2VFMADD132PD VFMADD132PS VFMADD132SD VFMADD132SS
3VFMADD213PD VFMADD213PS VFMADD213SD VFMADD213SS
4VFMADD231PD VFMADD231PS VFMADD231SD VFMADD231SS
5VFMADDSUB132PD VFMADDSUB132PS VFMADDSUB213PD
6VFMADDSUB213PS VFMADDSUB231PD VFMADDSUB231PS
7VFMSUB132PD VFMSUB132PS VFMSUB132SD VFMSUB132SS
8VFMSUB213PD VFMSUB213PS VFMSUB213SD VFMSUB213SS
9VFMSUB231PD VFMSUB231PS VFMSUB231SD VFMSUB231SS
10VFMSUBADD132PD VFMSUBADD132PS VFMSUBADD213PD
11VFMSUBADD213PS VFMSUBADD231PD VFMSUBADD231PS
12VFNMADD132PD VFNMADD132PS VFNMADD132SD
13VFNMADD132SS VFNMADD213PD VFNMADD213PS
14VFNMADD213SD VFNMADD213SS VFNMADD231PD
15VFNMADD231PS VFNMADD231SD VFNMADD231SS
16VFNMSUB132PD VFNMSUB132PS VFNMSUB132SD
17VFNMSUB132SS VFNMSUB213PD VFNMSUB213PS
18VFNMSUB213SD VFNMSUB213SS VFNMSUB231PD
19VFNMSUB231PS VFNMSUB231SD VFNMSUB231SS
20
21 ; IFMA instructions
22VPMADD52HUQ VPMADD52LUQ